4) A spring is compressed stored 1000 j of potential energy in it.Two identical masses of 1 kg each are on either of side the spring.What is the velocity of the two balls after the spring is released?

5) A 2Kg mass is pushed with 10N of force and has 1N of friction opposing its motion.If it starts from rest how fast is it moving after 2s?

6)A wheel on a robot has a radius of0.2m and is rotating at 3rad/s. How far did it roll during 4s?

7) A 100kg crate is pushed on a flat surface with how much force before it starts to move assuming us=0.8?
Answer #1

4) U = 2*0.5*m*v^2

==> v = sqrt(U/m)

= sqrt(1000/1)

= 31.62 m/s

5) impulse = (10-1)*2 = 18 N/s

impulse = chnage in momentum

m*v = 18

v = 18/2 = 9 m/s

6) v = r*w = 0.2*3 = 0.6 m/s

d = v*t = 0.6*4 = 2.4 m

7) F = mue*m*g = 0.8*100*9.8 = 784 N